OpenStudy (anonymous):

Find the number a such that the line x = a divides the region bounded by the curves x = y^2 − 1 and the y-axis into 2 regions with equal area. Give your answer correct to 3 decimal places.

OpenStudy (anonymous):

Draw a graph first

OpenStudy (anonymous):

 I know it looks like this

OpenStudy (anonymous):

But i really don't know where to go from here.

OpenStudy (danjs):

Find the area of the bound region from -1 to 0, and the integrate from -1 to 'a' and set that equal to half the area

OpenStudy (danjs):

x=a will be the vertical line that divides it into 2 equal areas

OpenStudy (anonymous):

Do you know how to find the area between the curve and the line?

OpenStudy (anonymous):

The top will be \(x=a\) and the bottom will be \(x=y^2-1\).

OpenStudy (anonymous):

The bounds of integration will be when the lines intersect

OpenStudy (anonymous):

Yea but don't we have split this one in two because if you just take the definite intergral you would get zero

OpenStudy (anonymous):

\[ a = y^2-1 \implies 0=y^2-(1+a) \implies 0=(y-\sqrt{1+a})(y+\sqrt{1+a}) \]This is different of squares.

OpenStudy (freckles):